OpenStudy (anonymous):

a = a is an example of which algebraic property? Symmetric Property Transitive Property Reflexive Property Commutative Property of Multiplication

OpenStudy (anonymous):

reflexive, google it!

OpenStudy (anonymous):

3(a + b) = 3a + 3b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

distributive

OpenStudy (anonymous):

Distributive property

OpenStudy (anonymous):

Because we will distribute 3 with a and b :)

OpenStudy (anonymous):

\[Solve =\frac{ x }{ 3 } = 12\]

OpenStudy (anonymous):

x=36 multiply both sides of equation by 3

OpenStudy (anonymous):

Yess :)

OpenStudy (anonymous):

solve \[solve \frac{ y }{ -2 } +5 = -4\]

OpenStudy (anonymous):

y=18 multiply both sides by -2 y-10=8 add 10 to both sides y=18

OpenStudy (anonymous):

Solve: −2(x − 3) = −6 0 −6 6 −12

OpenStudy (anonymous):

@Kerose Cool xD

OpenStudy (anonymous):

-2x+6=-6 -2x=-6-6 -2x=-12 x=-12/-2 x=6

OpenStudy (anonymous):

Solve: −4(x + 1) − 3 = −3(x − 4) −19 19 2 0

OpenStudy (anonymous):

-4x-4-3=-3x+12 -4x+3x=12+4+3 -x=19 x=-19

OpenStudy (anonymous):

Make different threads for different questions :)